Where is 2/6 and 26/2 located on the numebr line?
Answer by stanbon(75887)

Where is 2/6 and 26/2 located on the numebr line?
------
2/6 = 1/3 is between 0 and 1
----
26/2 = 13
----------------
Cheers,
Stan H.